OpenStudy (anonymous):

Help! The state of Wyoming has roughly the shape of a rectangle that is 3.6 x 102 miles long and 2.8 x 102 miles high. What is the approximate area of Wyoming? Hint: The area of a rectangle is product of its length and width. PLS! :3

OpenStudy (jeansweets):

3.6*102 = 367.2 and 2.8*102=285.6 the formula is L*W= so that would mean Wyoming is roughly 104872.32 miles long